# Diversifying the Expert Knowledge for Task-Agnostic Pruning in Sparse Mixture-of-Experts

### v1

@@ Abstract
By increasing model parameters but activating them sparsely when performing a task, the use of Mixture-of-Experts (MoE) architecture significantly improves the performance of Large Language Models (LLMs) without increasing the inference cost. However, the memory consumption due to the growing number of experts presents a challenge to the deployment of these models in many real world settings.
Our empirical study reveals that some experts encode redundant knowledge during pre-training. We thus propose a method of grouping and pruning similar experts to improve model’s parameter efficiency. We validate the effectiveness of our method by pruning two state-of-the-art MoE models, Mixtral-8x7B and Mixtral-8x22B. Evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks. To facilitate future research, we will release our code and the pruned MoE models.
@@ 1 Introduction
There are many recent studies on pruning MoE models. One line of work utilizes task-specific information to prune irrelevant experts. For example, Chen et al. [3] prune the less frequently visited experts based on experiments on a range of tasks. Chowdhury et al. [5] find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage. Li et al. [21] merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning. Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks. He et al. [11] explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop. In comparison, Lu et al. [27] enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance. We illustrate the difference among these works in fig. 1. Although pruning MoE models in task-agnostic settings is of great practical value, this area has not been fully explored.
@@
In this work, we explore how to prune MoE models in a task-agnostic fashion. Our study is motivated by the finding that, given the same input, many experts respond similarly, indicating that these experts likely encode similar knowledge, and thus are somewhat redundant. We propose a method to improve model parameter efficiency by pruning redundant experts in two stages. As shown in fig. 1, we first identify and group similar experts in the feature space. Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers. We validate the effectiveness of our method by pruning experts for two state of the art MoE models, Mixtral-8x7B and Mixtral-8x22B. Evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
@@ 3.1 Task Definition
Several prior works [27, 11, 21] have identified that removing some experts in different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks. The remaining question is which expert(s) should be pruned in the MoE layer.
Specifically, for a MoE LLM $\mathcal{M}(\cdot;\mathcal{F})$, which consists of multiple MoE layers =$\mathcal{F}=\{F^{l}(\cdot;\Theta^{l},W^{l},K)\}_{l=1}^{L}$, we want to prune $r$ experts from the total $N$. This leads to the expert modules $\Theta$ being reduced to $\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{r}}\}$, and the routing matrix $W$ being reduced to $\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{r}}\}$, where $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{r}\}$ is the set of indices to delete. We aim to find the pruned $\hat{\mathcal{M}}$ by deleting the optimal indices set $s^{l}$ in different MoE layers $l$ that satisfy the following optimization objective:
@@
where $\mathcal{L}$ is the loss function, $\hat{\mathcal{M}}(\cdot;\hat{\mathcal{F}})$ is the pruned LLM by deleting $r$ experts from each MoE layer $F$ in $\mathcal{F}$, and $\mathcal{D}$ is the dataset.
This task is particularly challenging as this is an NP-hard search problem with $(C^{r}_{N})^{L}$ possible solutions. In this work, we propose to estimate expert similarity in the feature space for greedy structured pruning within each MoE layer to address this problem.
@@ 3.2 Expert Similarity in MoE layer
Remark.
We notice that a concurrent work [26] also studies expert similarity, but focuses on the cosine similarity between expert weights. Our work examines the differences between experts using a more generalized metric based on the correlation between inputs and outputs, and obtains a different observation: even without considering the last layer, expert similarity does not gradually decrease with increasing depth of layers. For example, we observe higher pairwise expert similarity at the $11$-th layer compared to the $6$-th layer, which is cross-verified by both distance metrics. Additionally, the experts in the $26$-th layer also demonstrate a higher similarity with others compared with shallow layers. We provide the full evaluation results on visualizing the expert similarity of MoE layers in Mixtral-8x7B and Mixtral-8x22B in appendix A.
@@ 3.3 Discover and Merge Similar Experts on Feature Space
Then, we initialize the $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the distance matrix $D$.
Last, we leverage the graph partition algorithm to split $\mathcal{G}$ into $r$ subgraphs $\{\mathcal{G}_{i}\}_{i=1}^{r}$. The experts indexed by $\mathcal{V}_{i}$ of $\mathcal{G}_{i}$ shares the most similar knowledge with each other, and shows much difference with experts indexed by $\mathcal{V}_{t}$ of $\mathcal{G}_{t}$, $t\neq i$. To achieve this goal, we adopt a two-step strategy. At the first step, we partition the $\mathcal{V}$ of $\mathcal{G}$ into two subsets $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, where $\mathcal{V}_{1}$ has $r-1$ nodes and $\mathcal{V}_{2}$ has $G-r+1$ nodes. Then, we split $\mathcal{V}_{1}$ into $r-1$ subsets, $\mathcal{V}_{1,1},...,\mathcal{V}_{1,r-1}$. The splitting process can be formulated as follows,
$\min\sum_{i,j\in\mathcal{V}}x_{i}x_{j}D_{ij}-\sum_{i,j\in\mathcal{V}}x_{i}(1-x_{j})D_{ij}$ (5)
$-\sum_{i,j\in\mathcal{V}}(1-x_{i})(1-x_{j})D_{ij},$
@@
where $x_{i}=1$ if node $i$ in group $\mathcal{V}_{1}$, and $x_{i}=0$ otherwise. The first term is used to minimize the distance between nodes in $\mathcal{V}_{1}$, the second term is usd to maximize the distance of nodes between $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, and the third term is ued to maximize the distance of nodes in $\mathcal{V}_{2}$.
When pruning $G$ experts to $r$ experts, there are $r-1$ experts keeping unaffected in the entire process. Our graph partition strategy can preserve the most knowledge while diversifying the expert features in each MoE layers.
Merging Similar Experts with Routers. To preserve the different knowledge learned by different models clustered in the same group, we merge these experts as well as their routers on the weight space as follows,
$\hat{\theta}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}\theta_{\mathcal{V}_{n}(i)},$ (6)
$\hat{W}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$
@@ 3.4 Practical Implementation
While we have developed a general strategy to prune sparse MoEs, several practical questions arise during implementation:
Q1: How to decide $\mathcal{M}(\cdot)$ and perform task-agnostic pruning on sparse MoEs?
Intuitively, given the same input, similar experts should generate similar outputs. Thus, we propose to use the mean squared error on the outputs of two experts to evaluate the similarity between them.
We follow previous work [27] to select few samples from pre-trained dataset C4 [32] as the calibration dataset to compute the expert feature representation at the third step for similar experts discovery. The C4 dataset comprises multiple domains of contents, which can be a good surrogate to simulate the performance of the task-agnostic pruning.
However, with the limited size of the calibration dataset, the estimation of expert similarity based on the aforementioned expert representation could be inaccurate due to biased sampling. To improve the generalization ability at the discovery stage, we randomly mix up the token embedding for data augmentation and then use the augmented data for computing the expert representation and similarity. In practice, we iteratively update the expert similarity matrix by continuously shuffling and mixing up the data for full augmentation.
Q2: Is it possible to prune sparse MoEs without using any data?
Experts in the MoE layer of current models typically have the same structure; for instance, in Mixtral-8x7B, each expert is a two-layer FFN with SiLU activation. By training on different routed tokens, the experts usually learn different distributions during the training process. This results in a disparity between the learned feature spaces of different experts, which is also potentially reflected by their weights.
Based on this analysis, we can use the weights as the expert representation to discover similar experts for pruning. In our implementation, we consider two strategies: (1) directly using the vectorized weights, which does not account for the structure of experts, and (2) leveraging the local linearity of neural networks to compute the approximated weight matrix, namely the surrogate weights in our paper.
Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$. For input $x$, the output is computed as $f(x;\theta)=\theta_{2}(\sigma(\theta_{1}x)\cdot\theta_{3}x)$. Then, the vectorized weight is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$.
Q3: How do we determine the value of $\alpha$ for merging similar experts?
Two strategies are explored in our implementation to merge similar experts.
One strategy is to simply merge experts and gates with the same coefficients, i.e., $\alpha_{1}=\alpha_{2}=...=\alpha_{|\mathcal{V}_{n}|}=\frac{1}{|\mathcal{V}_{n}|}$, which encourages experts contributing equally.
The other strategy is to learn the merging coefficients for different experts towards minimizing the following loss function,
@@
$\hat{W}_{n}=\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$
where the ground-truth is the output of original MoE layer $y=F(x;\Theta,W,K)$, and we jointly optimize $\lambda$ and $\alpha$ for different merging groups.
@@ 4 Evaluation
Table 1: Results on pruning the Mixtral-8x7B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Augmented: prunable experts discover using augmented data and merging with uniform coefficients; 3) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq. 7. The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
@@
Ours Vectorized $\bf{\theta}$ 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
Augmented 56.2/47.8 72.7/57.2 52.1/42.7 68.1/54.0 85.6/83.2 32.8/28.6 68.6/66.5 62.3/54.3
Learn 56.0/48.0 73.1/57.0 52.4/43.3 68.2/54.6 86.4/83.3 31.4/28.5 69.3/67.1 62.4/54.5
@@ 4.1 Experiment Setup
Studied Models. We take pruning Mixtral-8x7B and Mixtral-8x22B [14] as the example. The Mixtral-8x7B has $32$ sparse MoE-involved layers, in each there are $8$ experts. The hidden dimension for token embedding is $4,096$, and the intermediate feature dimension in each FFN of MoEs is $14,336$. The Mixtral-8x22B is similar with Mixtral-8x7B, but with $56$ sparse MoE layers. It has the token embedding dimension of $6,144$ and the intermediate feature dimension of $16,384$. During the inference, each token will select $2$ experts in each MoE layer.
@@
Compared Baselines. We take three advanced MoE pruning methods [11, 21, 27] as our baseline for comparison. Among them, router-guided merging [21] is initially designed for task-specific MoE pruning, where we set the target dataset as the samples from the pre-training dataset. For task-agnostic methods, we select the frequency-based pruning method Expert Trimming [27], also the count-guided strategy and loss-based pruning method [27], namely the Enumerate in our paper. Under the task-agnostic pruning setting, we disable all the fine-tuning stage of all methods for fair comparison.
Our Approaches. We use four metrics to prune the model, including 1) Vectorized $\theta$: concact the vectorized model parameter $\theta$ together as the expert representation for grouping, and apply the uniform merging on the experts and routers. 2) Surrogate $\theta$: multiply the model parameters together as the weight for one-layer surrogate FFN and use it as the expert representation. The merging strategy is uniform merging. 3) Augmented: compute the expert representation based on the expert outputs on the augmented inputs. The merging strategy is uniform merging. 4) Learn: Use the same grouping strategy with Augmented but learn to merge for experts and routers.
Implementations. We randomly select $128$ samples from the pre-training dataset C4 [32] as the calibration dataset. For the prunable candidate group discovery, we use the mean square error to evaluate the similarity between experts. For the learnable merging, we initialize the merging coefficients $\alpha$ as the uniform and set $\lambda$ as $1$. In our experiments, we utilize four NVIDIA A100 GPUs, each with 80GB of memory.
4.2 Pruning the Mixtral Architecture
We present the results of pruning the Mixtral architecture, including Mixtral-8x7B and Mixtral-8x22B, from 8 experts to 6 and 4 experts in each layer, respectively. The results on the two models are depicted in table 1 and table 2 respectively.
Results on Mixtral-8x7B. We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average. Compared with count-guided strategy [11] which just drops the experts less visited, router-guided strategy [21] has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge. Besides, compared with count-guided and enumerate strategies [27] which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers.
Results on Mixtral-8x22B. With more parameters introduced in Mixtral-8x22B, the search for optimal experts to drop and merge becomes more challenging. Unlike the results on pruning Mixtral-8x7B, the strategy using the surrogate $\theta$ achieves state-of-the-art results, with only a minor performance drop of $2.8\%$ on average compared to the full model.
At the same time, it should be noted that the use of data doesn’t help boost the pruning performance as well, where the augmented and learning-based method has an average performance drop of $1.05\%$ and $1.0\%$, respectively. Also, the routing similarity-based pruning suffers model collapse under both two settings. We argue it is due to the heavy imbalance between the used data and parameter space, leading to the computation of expert similarity over-fitted on the small calibration dataset.
@@ 4.2 Pruning the Mixtral Architecture
Table 2: Results on pruning the Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$, 2) Augmented, and 3) Learn. The first column indicate the results on pruned model with $6$ experts, while the second column present the result of $4$ experts.
@@
Ours Vectorized $\bf{\theta}$ 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
Augmented 62.3/57.8 78.5/69.7 60.2/51.3 73.4/64.2 87.6/83.1 35.8/33.2 71.1/68.1 67.0/61.1
Learn 61.4/57.9 78.3/70.1 61.2/51.4 72.8/65.0 88.2/84.9 35.6/32.7 70.5/67.3 66.9/61.3

### v2

@@ Abstract
By increasing model parameters but activating them sparsely when performing a task, the use of Mixture-of-Experts (MoE) architecture significantly improves the performance of Large Language Models (LLMs) without increasing the inference cost. However, the memory consumption due to the growing number of experts presents a challenge to the deployment of these models in many real-world settings.
Our empirical study reveals that some experts encode redundant knowledge during pre-training. We thus propose a method of grouping and pruning similar experts to improve the model’s parameter efficiency. We validate the effectiveness of our method by pruning three state-of-the-art MoE architectures, including Mixtral, Deepseek-MoE, and Qwen. The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks. We will release our code to facilitate future research.
@@ 1 Introduction
There have been several studies on pruning MoE models. One line of work utilizes task-specific information to prune irrelevant experts. For example, Chen et al. [3] prune the less frequently visited experts based on experiments on a range of tasks. Chowdhury et al. [5] find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage. Li et al. [21] merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning. Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks. He et al. [11] explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop. In comparison, Lu et al. [26] enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance. We illustrate the difference among these works in fig. 1. Although pruning MoE models in task-agnostic settings is of great practical value, this area has not been fully explored.
@@
In this work, we explore how to prune MoE models in a task-agnostic fashion. Our study is motivated by the finding that, given the same input, many experts respond similarly, indicating that these experts likely encode similar knowledge, and thus are somewhat redundant. We propose a method to improve model parameter efficiency by pruning redundant experts in two stages. As shown in fig. 1, we first identify and group similar experts in the feature space. Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers. We validate the effectiveness of our method by pruning experts for three state-of-the-art MoE architectures, including Mixtral (Mixtral-8x7B and Mixtral-8x22B), Deepseek-MoE (Deepseek-MoE-16B) and Qwen (Qwen2-57B-14A). The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
@@ 3.1 Task Definition
Previous works [26, 11, 21] have identified that removing some experts from different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks. The remaining question is which expert(s) should be pruned in the MoE layer.
Specifically, for the $l$-th MoE layer $F^{l}(\cdot;\Theta^{l},W^{l},K)$ in the given LLM $\mathcal{M}(\cdot;\mathcal{F})$, we want to prune $r$ experts from the total of $N$, $l=1,2,...,L$. This leads to the expert modules $\Theta$ being reduced to $\hat{\Theta}^{l}=\Theta^{l}\setminus\{\theta^{l}_{s^{l}_{1}},\theta^{l}_{s^{l}_{2}},\ldots,\theta^{l}_{s^{l}_{r}}\}$, and the routing matrix $W$ being reduced to $\hat{W}^{l}=W^{l}\setminus\{W^{l}_{s^{l}_{1}},W^{l}_{s^{l}_{2}},\ldots,W^{l}_{s^{l}_{r}}\}$, where $s^{l}=\{s^{l}_{1},s^{l}_{2},\ldots,s^{l}_{r}\}$ is the set of indices to delete. We aim to find the pruned $\hat{\mathcal{M}}$ by deleting the optimal indices set $s^{l}$ in different MoE layers $l$ towards the following optimization objective:
@@
where $\mathcal{L}$ is the loss function, $\hat{\mathcal{M}}(\cdot;\hat{\mathcal{F}})$ is the new LLM with pruned MoE layer $\hat{\mathcal{F}}$, and $\mathcal{D}$ is any dataset under the task-agnostic setting.
This task is particularly challenging, as this is an NP-hard search problem with $(C^{r}_{N})^{L}$ possible solutions. In this work, we propose to estimate the expert similarity in the feature space for greedy structured pruning within each MoE layer to address this problem.
@@ 3.3 Discover and Merge Similar Experts on Feature Space
Then, we initialize $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the distance matrix $D$.
Last, we split $\mathcal{G}$ into $r$ subgraphs $\{\mathcal{G}_{i}\}_{i=1}^{r}$ to group similar experts. The experts indexed by $\mathcal{V}_{i}$ of $\mathcal{G}_{i}$ share the most similar knowledge with each other and show much difference with experts indexed by $\mathcal{V}_{t}$ of $\mathcal{G}_{t}$, $t\neq i$. This can be formulated as the follows,
$\min\sum_{i=1}^{r}\left(\sum_{j,k\in\mathcal{V}_{i}}D_{jk}-\sum_{t\neq i}^{r}\sum_{j\in\mathcal{V}_{i},k\in\mathcal{V}_{t}}D_{jk}\right),$ (5)
@@
where The objective function seeks to minimize the intra-group similarity (first term) and maximize the inter-group difference (second term) based on the pairwise similarity $D_{jk}$ based on expert $i$ and $j$.
Merging Similar Experts with Routers. To diversity the experts and preserve the different knowledge learned by different models clustered in the same group [37, 35], we merge the clusted experts with their routers on the weight space as follows,
$\hat{\theta}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}\theta_{\mathcal{V}_{n}(i)},~{}~{}\hat{W}_{n}$ $\leftarrow\sum_{i=1}^{|\mathcal{V}_{n}|}\alpha_{i}W_{\mathcal{V}_{n}(i)},$ (6)
@@ 3.4 Practical Implementation
On the discovery step. We propose two strategies for computing the disparity matrix $D$, including the data-centric strategy and model-centric strategy.
For data-centric strategy, we follow previous work [26] to use the pre-trained dataset C4, which can be a good surrogate to simulate the performance of the task-agnostic pruning. To improve the generalization ability at the discovery stage, we randomly mix the token embedding for data augmentation and then use the augmented data to calculate the expert output as the expert representation.
For model-centric strategy, we propose two ways to prune models with only expert weights, which encode the dataset information during the training process and can be a good agency for expert representation. The first way is to leverage the vectorized weights directly. The other way is to leverage the local linearity of neural networks to compute the surrogate weight matrix.
Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$. For input $x$, the output is computed as $f(x;\theta)=\theta_{2}(\sigma(\theta_{1}x)\cdot\theta_{3}x)$. Then, the vectorized weight is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$. Compared to vectorized weights, surrogate weights are more flexible with model size, showing more stable performance in pruning models (see section A.2).
On the merging step. We consider three approaches in practice deciding $\alpha$ to merge similar grouped experts. The first is to only maintain the experts with the maximum visiting frequency and drop all the others. The second is to set $\alpha_{i}=\frac{1}{|\mathcal{V}_{n}|}$, which uniformly assembles all the experts grouped. The last one is learning $\alpha$ to merge the grouped experts by minimizing the following loss function,
@@
where the ground-truth is the output of original MoE layer $y=F(x;\Theta,W,K)$, and we jointly optimize $\lambda$ and $\alpha$ for different merging groups. Among these three strategies, we empirically find that the uniform souping strategy offers more stable performance and greater efficiency in model pruning (see section A.2 ). While the learning strategy can yield slightly better performance, it is more time-consuming due to the need for tuning the parameter $\alpha$.
@@ 4 Evaluation
Table 1: Results on pruning the Mixtral-8x7B from 8 experts to 6 and 4 experts in each MoE layer. The first and second columns respectively indicate the results of the pruned model with $6$ and $4$ experts.
@@
Ours Model-centric 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
Data-centric 56.0/48.0 73.1/57.0 52.4/43.3 68.2/54.6 86.4/83.3 31.4/28.5 69.3/67.1 62.4/54.5
@@ 4.1 Experiment Setup
Studied Models. We take pruning three MoE-based architectures as an example, including the Mixtral, Deepseek, and Qwen. For the Mixtral architecture, the Mixtral-8x7B has $32$ sparse MoE-involved layers, in each there are $8$ experts. The Mixtral-8x22B is similar to Mixtral-8x7B but with $56$ sparse MoE layers. During the inference, each token will select $2$ experts in each MoE layer. The deepseek model has $28$ layers, and there are $64$ experts in each layer. Each token will pass $2$ shared experts and select $6$ experts during the inference. The Qwen model also has $28$ MoE layers with $64$ experts in each layer but will activate $8$ experts during the inference.
@@
Pruning Methods. We take three advanced MoE pruning methods [11, 21, 26] as our baseline for comparison. Among them, router-guided merging [21] is initially designed for task-specific MoE pruning, where we set the target dataset as the samples from the pre-training dataset. For task-agnostic methods, we select the frequency-based pruning method Expert Trimming [26], also the count-guided strategy and loss-based pruning method [26], namely the Enumerate in our paper. Under the task-agnostic pruning setting, we disable all the fine-tuning stage of all methods for fair comparison.
For our method, we respectively report the best results of our data-centric and model-centric methods in pruning models. Following the previous setting [26], we use $128$ samples in C4 for computation in data-centric pruning methods, while the model-centric method doesn’t rely on the data. More detailed results on different implementations of discovery and merging steps (e.g., vectorized and surrogate weight strategy at the discovery step, max or learning strategy at the merging step) are provided in section A.2.
@@
Table 2: Results on pruning the Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer. The first and second columns respectively indicate the results of the pruned model with $6$ and $4$ experts.
@@
Ours Model-centric 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
Data-centric 62.3/57.8 78.5/69.7 60.2/51.3 73.4/64.2 87.6/83.1 35.8/33.2 71.1/68.1 67.0/61.1
4.2 Pruning the Mixtral Architecture
We present the results of the pruning of the Mixtral architecture, including Mixtral-8x7B and Mixtral-8x22B. Both have $8$ experts in each MoE layer. We apply different methods to prune them from 8 experts to 6 and 4 experts in each layer, respectively. The results of the two models are shown in table 1 and table 2 respectively.
Results on Mixtral-8x7B. We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average. Compared with count-guided strategy [11] which just drops the experts less visited, router-guided strategy [21] has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge. Besides, compared with count-guided and enumerate strategies [26] which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers. We can also notice that the model-centric method surpasses all the other data-involved pruning baseline methods. This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning.
Results on Mixtral-8x22B. In this experiment, the model-centric method achieves the best result with only a minor performance drop of $2.8\%$ on average compared to the full model. Our proposed data-centric method achieves the second to the last performance, with a gap of $0.8\%$ performance gap compared with the runner-up method. It should be noted that the routing similarity-based pruning suffers model collapse under both two settings, showing a less robustness compared with other methods and metrics for similarity computation.
4.3 Pruning the Deepseek Model
We also evaluate the effectiveness of our proposed pruning method on compressing the Deepseek architecture. Specifically, we pruned the experts of Deepseek-MoE-16B, reducing the number of experts from 64 to 48 using different model pruning strategies. The results are illustrated in fig. 4.
Notably, even after pruning one-third of the experts in Deepseek-MoE-16B, our data-centric strategy maintains an impressive average performance of $50.9\%$, with only a $3.1\%$ reduction in performance compared to the full model. In the evaluation of the most challenging MMLU task, our model-centric strategy demonstrates superior performance in most cases, particularly when reducing the number of experts from $62$ to $52$. It consistently outperforms the runner-up baseline method, achieving a clear performance advantage of $2.6\%$.
Figure 4: Results on pruning the Deepseek-MoE-16B with different strategies.
4.4 Pruning the Qwen Model
Figure 5: Results on pruning the Qwen2-57B-14A with different strategies.
The Qwen architecture is also utilized in our experiments. We study the Qwen2-57B-14A to evaluate the performance of different model pruning strategies. We prune the experts from $64$ experts to $48$ experts in each layer and evaluate the performance on different tasks. The results are reported in fig. 5.
When pruning $\frac{1}{3}$ experts of the full model, the data-centric strategy delivers the best performance, outperforming both our model-centric method and the count-guided baseline method. On the MMLU task, our model-centric strategy achieves approximately $1.5\%$ better performance compared to the count-guided pruning strategy.